The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

                                 mcnews@astro.uiuc.edu
                 http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
                         http://www.astro.uni-bonn.de/~mcnews/

No. 60 November 11, 2001

Contents

Abstracts of 15 refereed papers  1
Abstracts of 2 non-refereed papers 10
Job opportunity  11

Abstracts of Refereed Papers

The Hi Environment of Three Superbubbles in the Large Magellanic Cloud

M. S. Oey^{1}, B. Groves^{2}, L. Staveley-Smith^{3}, and R. C. Smith^{4}

1 Lowell Observatory, 1400 W. Mars Hill Road, Flagstaff, AZ 86001, USA 2 Research School of Astronomy and Astrophysics, ANU, Private Bag, Weston Creek P.O., ACT 2611, Australia 3 Australia Telescope National Facility, CSIRO, P.O. Box 76, Epping, NSW 1710, Australia 4 Cerro Tololo Inter-American Observatory, Casilla 603, La Serena, Chile

The ambient interstellar environment of wind- and supernova-driven superbubbles strongly affects their evolution, but its properties are rarely well-determined. We have therefore obtained H i aperture synthesis imaging of the environment around three similar, optically-selected superbubble nebulae in the Large Magellanic Cloud. The resulting H i maps show that the ambient gas distribution around these superbubbles differ to an extreme: DEM L25 shows no neutral shell component, but is nestled within an H i hole; DEM L50 shows a massive neutral shell component, but is otherwise within an H i void; and DEM L301 shows no correspondence at all between the optical nebula and H i distribution. There is also poor correspondence between the H i and optical kinematics. These results strongly caution against inferring properties of the ambient neutral environment of individual superbubbles without direct observations. Finally, all three objects show some evidence of shock activity.

Accepted by: The Astronomical Journal

For preprints, contact Sally.Oey@Lowell.edu
Also available from the URL http://www.lowell.edu/~oey

Monitoring the Evolution of the X-ray Remnant of SN 1987A

Sangwook Park^{1}, David N. Burrows^{1}, Gordon P. Garmire^{1}, John A. Nousek^{1}, Richard McCray^{2}, Eli Michael^{2}, and Svetozar Zhekov^{2}

1 Dept. of Astronomy & Astrophysics, Pennsylvania State University, 525 Davey Lab., University Park, PA 16802, USA 2 Joint Institute for Laboratory Astrophysics, University of Colorado, Campus Box 440, Boulder, CO 80309-0440, USA

We report on the results of our monitoring program of the remnant of SN 1987A with the Advanced CCD Imaging Spectrometer (ACIS) on board the Chandra X-ray Observatory. Two new observations have been performed in AO2, bringing the total to four monitoring observations over the past two years. Over this time period, new techniques for correction of “Charge Transfer Inefficiency (CTI)” and for use of charge spreading to provide angular resolution somewhat better than the pixel size of the CCD detector have become available at Penn State. We have processed all four observations using sub-pixel resolution to obtain the highest possible angular resolution, and using our CTI correction software to provide more reliable spectral analysis and flux estimations.

The high angular resolution images indicate that the X-ray bright knots are convincingly correlated with the optical spots, primarily at ∼ <1 keV, while higher energy photons are very well correlated with radio images. Our data also provide marginal evidence for radial expansion of the X-ray remnant at a rate of 5200 ± 2100 km s^{−1}. The X-ray flux appears to linearly increase by ∼60% over the 18 month period of these observations. The spectrum is dominated by broad complexes of atomic emission lines and can be fit with a simple model of a plane-parallel shock with electron temperatures of kT ∼ 2 − 4 keV and a postshock electron density of n_{e} ∼ 210 − 420 cm^{−3}. The implied 0.5 − 10 keV band luminosity in 2001 April is ∼1.3 × 10^{35} ergs s^{−1}; as of that date, we still observe no direct evidence for the central point source, with an upper limit on the observed luminosity of L_{X} ∼ 5.5 × 10^{33} ergs s^{−1} in the 2 − 10 keV band.

Accepted by: The Astrophysical Journal For preprints, contact park@astro.psu.edu

STIS spectroscopy of newborn massive stars in SMC N81

M. Heydari-Malayeri^{1}, M. R. Rosa^{2}, D. Schaerer^{3}, F. Martins^{3}, and V. Charmandaris^{4}

1 Observatoire de Paris, 61 Avenue de l’Observatoire, F-75014 Paris, France 2 Space Telescope European Coordinating Facility, ESO, Karl-Schwarzschild-Str. 2, D-85748 Garching bei München, Germany 3 Laboratoire d’Astrophysique, Observatoire Midi-Pyrénées, 14, Avenue E. Belin, F-31400 Toulouse, France 4 Cornell University, Astronomy Department, 106 Space Sciences Bldg., Ithaca, NY 14853, USA

Using Hubble Space Telescope observations with STIS, we study the main exciting stars of N81, a high excitation compact H ii region in the Small Magellanic Cloud (SMC). These far UV observations are the first spectroscopic measurements of stars in such a region and reveal features characteristic of an O6–O8 stellar type. The astonishing weakness of their wind profiles and their sub-luminosity (up to ∼ 2 mag fainter in M_{V} than the corresponding dwarfs) make these stars a unique stellar population in the Magellanic Clouds. Our analysis suggests that they are probably in the Hertzsprung-Russell diagram locus of a particularly young class of massive stars, the so-called Vz luminosity class, as they are arriving on the zero age main sequence.

Accepted by: Astronomy & Astrophysics For preprints, contact heydari@obspm.fr

Also available from the URL http://wwwusr.obspm.fr/∼heydari/projects/N81/stis.ps.gz or:
http://xxx.lanl.gov/ps/astro-ph/0110576

Editors’ note: A press release accompanying this paper can be found at

   http://sci.esa.int/hubble/news/index.cfm?oid=28750

Resolving the compact H ii regions in N160A with HST

M. Heydari-Malayeri^{1}, V. Charmandaris^{2}, L. Deharveng^{3}, F. Meynadier^{1}, M.R. Rosa^{4}, D. Schaerer^{5}, & H. Zinnecker^{6}

1Observatoire de Paris, DEMIRM, 61 Ave de l’Observatoire, F–75014 Paris, France 2Astronomy Department, Cornell Univ., Ithaca, NY 14853, USA 3Observatoire de Marseille, 2 Place Le Verrier, F-13248 Marseille Cedex 4, France 4STECF - European Southern Observatory, D-85748 Garching, Germany 5Observatoire Midi-Pyrenees, 14, Avenue E. Belin, F-31400 Toulouse, France 6Astrophysikalisches Institut Potsdam, D-14482 Potsdam, Germany

Using high-resolution imaging with the Hubble Space Telescope, we study the Large Magellanic Cloud H ii region N160A and uncover several striking features of this complex massive star-forming site. The two compact high excitation H ii blobs (HEBs) A1 and A2 are for the first time resolved and their stellar content and morphology is revealed. A1, being of higher excitation, is powered by a single massive star whose strong wind has created a surrounding bubble. A2 harbors several exciting stars enshrouded inside large quantities of dust. The whole N160A nebula is energized by three star clusters for which we obtain photometry and study their color-magnitude diagram. The H ii region is particularly dusty, with extinction values reaching an A_{V} ∼ 2.5 mag in the visible, and it is separated from the molecular cloud by an outstanding ionization front. A previously detected infrared young stellar object is also accurately located with respect to the H ii region.

Accepted by: Astronomy and Astrophysics

For preprints, contact mohammad.heydari-malayeri@obspm.fr
Also available from the URL http://arXiv.org/abs/astro-ph/0110622

Mass segregation in young compact star clusters in the Large Magellanic Cloud: I. Data and Luminosity Functions

R. de Grijs^{1}, R.A. Johnson^{1,2}, G.F. Gilmore^{1} and C.M. Frayn^{1}

1 Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 0HA, UK 2 European Southern Observatory, Alonso de Cordova 3107, Santiago 19, Chile

We have undertaken a detailed analysis of HST/WFPC2 and STIS imaging observations, and of supplementary wide-field ground-based observations obtained with the NTT of two young (∼ 10 − 25 Myr) compact star clusters in the LMC, NGC 1805 and NGC 1818. The ultimate goal of our work is to improve our understanding of the degree of primordial mass segregation in star clusters. This is crucial for the interpretation of observational luminosity functions (LFs) in terms of the initial mass function (IMF), and for constraining the universality of the IMF.

We present evidence for strong luminosity segregation in both clusters. The LF slopes steepen with cluster radius; in both NGC 1805 and NGC 1818 the LF slopes reach a stable level well beyond the clusters’ core or half-light radii. In addition, the brightest cluster stars are strongly concentrated within the inner ∼ 4R_{hl}.

The global cluster LF, although strongly nonlinear, is fairly well approximated by the core or half-light LF; the (annular) LFs at these radii are dominated by the segregated high-luminosity stars, however.

We present tentative evidence for the presence of an excess number of bright stars surrounding NGC 1818, for which we argue that they are most likely massive stars that have been collisionally ejected from the cluster core. We therefore suggest that the cores of massive young stars clusters undergo significant dynamical evolution, even on time-scales as short as ∼ 25 Myr.

Submitted to: Monthly Notices of the Royal Astronomical Society For preprints, contact grijs@ast.cam.ac.uk

Mass segregation in young compact star clusters in the Large Magellanic Cloud: II. Mass Functions

R. de Grijs^{1}, G.F. Gilmore^{1}, R.A. Johnson^{1,2}, and A.D. Mackey^{1}

1 Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 0HA, UK 2 European Southern Observatory, Alonso de Cordova 3107, Santiago 19, Chile

We review the complications involved in the conversion of stellar luminosities into masses and apply a range of mass-to-luminosity relations to our Hubble Space Telescope observations of the young LMC star clusters NGC 1805 and NGC 1818.

Both the radial dependence of the mass function (MF) and the dependence of the cluster core radii on mass indicate clear mass segregation in both clusters at radii r ≤ 20 − 30′′, for masses in excess of ∼ 1.6 − 2.5M_{⊙}. This result does not depend on the mass range used to fit the slopes or the metallicity assumed. It is clear that the cluster MFs, at any radius, are not simple power laws.

The global and the annular MFs near the core radii appear to be consistent with the Salpeter IMF slope in the mass range (−0.15 ≤ log m/M_{⊙} ≤ 0.85), the MFs beyond r ≥ 30′′ have significantly steeper slopes.

We estimate that while the NGC 1818 cluster core is between ∼ 5 and ∼ 30 crossing times old, the core of NGC 1805 is likely ≤ 3 − 4 crossing times old. However, since strong mass segregation is observed out to ∼ 6R_{core} and ∼ 3R_{core} in NGC 1805 and NGC 1818, respectively, it is most likely that significant primordial mass segregation was present in both clusters, particularly in NGC 1805.

Submitted to: Monthly Notices of the Royal Astronomical Society For preprints, contact grijs@ast.cam.ac.uk

An empirical method to estimate the LMC distance using B-stars in eclipsing binary systems

M. Salaris^{1,2} and M.A.T. Groenewegen^{3,4}

1 Astrophysics Research Institute, Liverpool John Moores University, Twelve Quays House, Egerton Wharf, Birkenhead CH41 1LD, UK 2 Max-Planck-Institut für Astrophysik, Karl-Schwarzschild-Straße 1, D-85748 Garching, Germany 3 Instituut voor Sterrenkunde, PACS-ICC, Celestijnenlaan 200B, B-3001 Heverlee, Belgium 4 European Southern Observatory, EIS-team, Karl-Schwarzschild-Straße 2, D-85740 Garching, Germany

We present a new method to determine the distance to B-stars in eclipsing binary systems. The method is completely empirical, and it is based on the existence of a very tight linear relationship between the V -band “zero magnitude angular diameter” and the Strömgren colour index c_{1} for B-stars; we have empirically calibrated this relationship using local single B-stars with accurate angular diameters, and B-stars in eclipsing binaries with precise radii and parallax determinations. By studying the differential behaviour of this relationship as predicted by theoretical stellar evolution models, we find that it is independent of the stellar metallicities for a range of [Fe/H] values between the solar one and that of young stars in the Magellanic Clouds. The method, which also provides the value of the reddening to the system, is discussed in detail, together with a thorough estimate of the associated errors. We conclude that accurate Strömgren photometry obtainable with 1.5 m-class telescopes of the LMC eclipsing binaries HV 2274 and HV 982 will allow to obtain an empirical LMC distance with an accuracy of the order of 0.13 mag.

Accepted by: Astronomy & Astrophysics

For preprints, contact ms@astro.livjm.ac.uk
Also available from the URL http://arXiv.org/abs/astro-ph/0111082

Period Changes of the LMC Cepheids determined from the Harvard and OGLE/ASAS data

P. Pietrukowicz^{1}

1 Warsaw University Observatoru, Al. Ujazdowskie 4, 00-478, Poland

Observations of Cepheids in the Large Magellanic Cloud, made over last several decades, allow us to search for evolutionary period changes. None of the Cepheid from our sample of 378 stars stopped pulsating. Also none of them showed a large period change which could indicate mode switching. However for Cepheids with log P > 0.9 we found significant period changes, positive as well as negative. A comparison between the observed period changes and theoretical predictions shows moderate agreement with some models (Bono et al. 2000), and a very large disagreement with others (ABHA - Alibert, Baraffe, Hauschildt, Allard 1999). The large differences between the models are likely caused by the very high sensitivity of stellar evolution during core helium burning phase to even small changes in the input physics, as discovered by Lauterborn, Refsdal and Weigert (1971).

Accepted by: Acta Astronomica

For preprints, contact pietruk@astrouw.edu.pl
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0108407

Globular Clusters and the Mira Period-Luminosity Relation

Michael Feast^{1}, Patricia Whitelock^{2} and John Menzies^{2}

1 Astronomy Dept. University of Cape Town, Rondebosch, 7701, South Africa 2 South African Astronomical Observatory, P.O. Box 9, Observatory, 7935, South Africa

A globular cluster distance scale based on Hipparcos parallaxes of subdwarfs has been used to derive estimates of M_{K} for cluster Miras, including one in the SMC globular cluster NGC121. These lead to a zero-point of the infrared period-luminosity relation, PL(K), in good agreement with that derived from Hipparcos parallaxes of nearby field Miras. The mean of these two estimates together with data on LMC Miras yields an LMC distance modulus of 18.60 ± 0.10 in evident agreement with a metallicity corrected Cepheid modulus (18.59± ∼ 0.10).

The use of luminous AGB stars as extragalactic population indicators is also discussed.

Accepted by: Monthly Notices of the Royal Astronomical Society

For preprints, contact mwf@artemisia.ast.uct.ac.za
Also available from the URL http://arXiv.org/abs/astro-ph/0111108

A UBVR CCD Survey of the Magellanic Clouds

P. Massey^{1}

1 Lowell Observatory, 1400 W. Mars Hill Road, Flagstaff, AZ 86001, USA

We present photometry and a preliminary interpretation of a UBVR survey of the Large and Small Magellanic Clouds, which covers 14.5 deg^{2} and 7.2 deg^{2}, respectively. This study is aimed at obtaining well-calibrated data on the brighter, massive stars, complementing recent, deeper CCD surveys. Our catalog contains 179,655 LMC and 84,995 SMC stars brighter than V ∼ 18.0, and is photometrically complete to U ∼ B ∼ V ∼ 15.7, and R ∼ 15.2, although stars in crowded regions are selectively missed. We provide tentative cross-reference between our catalog stars and the stars with existing spectroscopy. Our photometry agrees well with the photoelectric work in V and B–V, and agrees well for U–B for the bluest stars, but we find a large discrepancy (0.3 mag) in the U–B color at U–B∼ 0.0. Examination of the colors of stars with known spectral types suggests that the problem may lie with the photoelectric data. We discuss the population of stars seen towards the two Clouds, identifying the features in the color-magnitude diagram, and using existing spectroscopy to help construct H-R diagrams. We derive improved values for the blue to red star ratios in the two Clouds, emphasizing the uncertainties involved in this before additional spectroscopy. We compare the relative number of RSGs and Wolf-Rayet stars in the LMC and SMC with that of other galaxies in the Local Group, demonstrating a very strong, tight trend with metallicity, with the ratio changing by a factor of 160 from the SMC to M31. We also reinvestigate the initial mass function of the massive stars found outside of the OB associations. With the newer data, we find that the IMF slope of this field population is very steep, with Γ ∼ −4 ± 0.5, in agreement with our earlier work. This is in sharp contrast to the IMF slope found for the massive stars with OB associations (Γ ∼ −1.3). Although much more spectroscopy is needed to make this result firm, incompleteness can no longer be invoked as an explanation.

Accepted by: The Astrophysical Journal Supplement Series

Available from the URL ftp://ftp.lowell.edu/pub/massey/mcatlas.ps.gz

The Magellanic Clouds Photometric Survey:

The Small Magellanic Cloud Stellar Catalog and Extinction Map

Dennis Zaritsky^{1}, Jason Harris^{2}, Ian B. Thompson^{3}, Eva K. Grebel^{4}, and Philip Massey^{5}

1Steward Observatory, University of Arizona, Tucson, AZ 85721, USA 2Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 3Carnegie Observatories, 813 Santa Barbara St., Pasadena, CA 91101, USA 4Max Planck Institute for Astronomy, Königstuhl 17, D-69117 Heidelberg, Germany 5Lowell Observatory, 1400 West Mars Hill, Flagstaff, AZ 86001, USA

We present our catalog of U, B, V, and I stellar photometry of the central 18 sq. degree area of the Small Magellanic Cloud. We combine our data with the 2MASS and DENIS catalogs to provide, when available, U through K_{S} data for stars. Internal and external astrometric and photometric tests using existing optical photometry (U, B, and V from Massey’s bright star catalog; B, V, and I from the microlensing database of OGLE and I from the near-infrared sky survey DENIS) are used to determine the observational uncertainties and identify systematic errors. We fit stellar atmosphere models to the optical data to check the consistency of the photometry for individual stars across the passbands and to estimate the line-of-sight extinction. Finally, we use the estimated line-of-sight extinctions to produce an extinction map across the Small Magellanic Cloud and investigate the nature of extinction as a function of stellar population.

Accepted by: The Astronomical Journal

Also available from the URL http://arXiv.org/abs/astro-ph/0110665

The recent structural evolution of the SMC

F. Maragoudaki^{1}, M. Kontizas^{1}, D.H. Morgan^{2}, E. Kontizas^{3}, A. Dapergolas^{3}, and E. Livanou^{1}

1 Section of Astrophysics Astronomy & Mechanics, Dpt. of Physics, University of Athens, GR-157 83 Athens, Greece 2 Institute for Astronomy, University of Edinburgh, Blackford Hill, Edinburgh, EH9 3HJ, UK 3 Institute of Astronomy and Astrophysics, National Observatory of Athens, P.O. Box 20048, GR-118 10 Athens, Greece

We investigate the spatial distribution of stars of different ages towards the main body of the SMC, in an attempt to further understand the nature of the complex structure of the SMC. The old stellar population of the galaxy shows a rather regular and smooth distribution which is typical for a spheroidal body. On the contrary, the distribution of the younger stellar component is highly asymmetric and irregular giving evidence for the severe impact of the SMC during its close encounter with the LMC some 0.2 to 0.4 Gyr ago. In a series of isodensity contour maps of stars within selected ages, we show the recent structural evolution of the SMC.

Accepted by: Astronomy & Asrophysics For preprints, contact M. Kontizas, mkontiza@cc.uoa.gr

The Optical Gravitational Lensing Experiment. Difference Image Analysis of LMC and SMC data. The Method

K. Żebruń^{1,2}, I. Soszyński^{1,2} and P. R. Woźniak^{3,2}

1 Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warsaw, Poland 2 Princeton University Observatory, Princeton, NJ 08544–1001, USA 3 Los Alamos National Observatory, MS-D436, Los Alamos NM 85745, USA

We describe the Difference Image Analysis (DIA) algorithms and software used to analyze four years (1997-2000) of OGLE-II photometric monitoring of the Magellanic Clouds, the calibration, the photometric error analysis, and the search for variable stars. A preliminary analysis of photometric errors is based on the field LMC SC2. A full catalog of about 70,000 variable stars will be presented in a separate publication.

Submitted to: Acta Astronomica

For preprints, contact zebrun@astrouw.edu.pl
Also available from the URL http://arXiv.org/abs/astro-ph/0110612

The Optical Gravitational Lensing Experiment. Difference Image Analysis of LMC and SMC data. The Catalog

K. Żebruń^{1,2}, I. Soszyński^{1,2}, P. R. Woźniak^{3,2}, A. Udalski^{1}, M. Kubiak^{1}, M. Szymański^{1}, G Pietrzyński^{1,4}, O. Szewczyk^{1} and Ł. Wyrzykowski^{1}.

1 Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warsaw, Poland 2 Princeton University Observatory, Princeton, NJ 08544–1001, USA 3 Los Alamos National Observatory, MS-D436, Los Alamos NM 85745, USA 4 Universidad de Concepción, Departamento de Fisica, Casilla 160–C, Concepción, Chile

We present the first edition of a catalog of variable stars found in the Magellanic Clouds using OGLE-II data obtained during four years: 1997-2000. The catalog covers about 7 square degrees of the sky – 21 fields in the Large Magellanic Cloud and 11 fields in the Small Magellanic Cloud. All variables were found with the Difference Image Analysis (DIA) software. The catalog is divided into two sections. The DC section contains FITS reference images (obtained by co-adding 20 best frames for each field) and profile photometry (DoPHOT) of all variable stars on those images. The AC section contains flux variations and magnitudes of detected variable stars obtained with DIA as well as with DoPHOT. The results are available in electronic form via FTP and through WWW interface from the OGLE internet archive. Typically, there are about 400 I-band data points and about 30 V and B-band data points for over 68,000 variables. A detailed analysis and classification of variable stars will be presented elsewhere. The FTP catalog contains approximately 2Gbytes of data.

Submitted to: Acta Astronomica

For preprints, contact zebrun@astrouw.edu.pl
Also available from the URL http://arXiv.org/abs/astro-ph/0110623

The Galactic Halo from microlensing

A. Milsztajn^{1}

1 DSM-DAPNIA-Service de Physique des Particules, CEA-Saclay, F-91191 Gif-sur-Yvette, France

The present status of the search for galactic dark matter through microlensing of Magellanic Clouds stars is reviewed. After recalling the basics of microlensing, I describe the candidate microlensed stars towards the LMC and the SMC. I then assess the results, with emphasis on the unresolved issue of where the lensing objects are located. I mention possible ways to solve this question in the near future.

Accepted by: Space Science Reviews (Proceedings of the ISSI workshop “Matter in the Universe”, Bern, March 2001) For preprints, contact amilsztajn@cea.fr

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0111117,
or (soon) from the URL http://www-dapnia.cea.fr/Doc/Publications/Spp/spp01.php
as document DAPNIA/SPP-01-15

Abstracts of Non-Refereed Papers

CMD’s, RR Lyrae’s, clump stars and reddening in the LMC

G. Clementini^{1}, A. Bragaglia^{1}, M. Maio^{1}, E. Carretta^{2}, R.G. Gratton^{2} and L. Di Fabrizio^{3}

1 Osservatorio Astronomico di Bologna, Via Ranzani n.1, 40127 Bologna, Italy 2 Osservatorio Astronomico di Padova, Vicolo dell’Osservatorio n. 5, 35122 Padova, Italy 3 Centro Galileo Galilei & Telescopio Nazionale Galileo, PO Box 565, 38700 S. Cruz de La Palma, Spain

We present new B, V, I photometry in the standard Johnson-Cousins photometric system of two regions near the bar of the Large Magellanic Cloud (LMC), one of which is very close to LMC field used for the ”Coimbra test on the synthetic CMD techniques”. The two areas contain 128 RR Lyrae variables and more than 8000 clump stars. Once combined with Clementini et al. 2001 (astro-ph/0007471) photometry the new data-set allows us to get (i) accurate CMD’s down to V ∼23 mag, (ii) full coverage of the B and V light variation for more than 80% of the RR Lyrae’s, (iii) a very precise estimate of the LMC RR Lyrae’s average apparent luminosity, (iv) an accurate estimate of the I luminosity of the LMC clump stars, and (v) an independent evaluation of the reddening in these regions of the LMC using the pulsational properties of the RR Lyrae’s in our sample. A detailed comparison is made between our, OGLE II and MACHO photometries. We discuss the impact of our new results on the distance to the LMC and on the short and long distance scale controversy.

To appear in: “Observed HR diagrams and stellar evolution: the interplay between observational constraints and theory” (Coimbra, June, 2001), ASP Conf. Ser., 2001, Thibault Lejeune and Joao Fernandes, eds. For preprints, contact gisella@bo.astro.it

Interstellar Media in the Magellanic Clouds and other Local Group Dwarf Galaxies

Eva K. Grebel

Max Planck Institute for Astronomy, Königstuhl 17, D-69117 Heidelberg, Germany

I review the properties of the interstellar medium in the Magellanic Clouds and Local Group dwarf galaxies. The more massive, star-forming galaxies show a complex, multi-phase ISM full of shells and holes ranging from very cold phases (a few 10 K) to extremely hot gas (> 10^{6} K). In environments with high UV radiation fields the formation of molecular gas is suppressed, while in dwarfs with low UV fields molecular gas can form at lower than typical Galactic column densities. There is evidence of ongoing interactions, gas accretion, and stripping of gas in the Local Group dwarfs. Some dwarf galaxies appear to be in various stages of transition from gas-rich to gas-poor systems. No ISM was found in the least massive dwarfs.

To appear in: Proceedings of the 17th IAP Colloquium on Gaseous Matter In Galaxies And Intergalactic Space, ed. R. Ferlet (Frontier Group: Paris)

Also available from the URL http://arXiv.org/abs/astro-ph/0110471

Job Opportunity

Postdoctoral position on: Massive star forming regions and starburst galaxies

The Geneva Observatory in Geneva, Switzerland, announces the availability of a research position at the postdoctoral level, open to applicants of all nationalities.

The succesfull candidate will work on a project aimed at studying the stellar populations and the interstellar medium of massive star forming regions from the local Universe to high redshift, involving both multi-wavelength observations and theoretical modeling. The successfull applicant will in particular have access to ground-based observational data covering the optical to mid-IR (including ESO/VLT data), and state-of-the-art modeling tools. He/she will mostly work in collaboration with Prof. Daniel Schaerer in Geneva.

The Geneva Observatory carries out observational, interpretative and theoretical research in fields including stellar evolution, stellar physics, galaxy evolution and dynamics, high energy astrophysics, and extra-solar planets.

The initial appointment will be for one to two years starting in fall 2002 (1 october 2002 or later). It is renewable.

Qualified candidates are encouraged to send their application including a CV, description of research experience and interests, and contact information of three references preferably via e-mail to Daniel Schaerer (schaerer@ast.obs-mip.fr), Laboratoire d’Astrophysique, 14, Av. E. Belin, F-31400 Toulouse, France.

All applications received by January 4, 2002 will receive full consideration. Informal enquiries with Daniel Schaerer (schaerer@ast.obs-mip.fr) are welcome.

PhD position on: Massive star forming regions and starburst galaxies

One or two PhD student positions at the Geneva Observatory are open on topics related to ”Massive Star Forming Regions from the local Universe to high redshift”, i.e. objects ranging from giant HII regions and starburst galaxies to the first galaxies (so-called Population III objects). The student will work with Prof. D. Schaerer and collaborators of the group. Observational and/or theoretical subjects are proposed.

Start of PhD: fall 2002 (1 october 2002 or later) For more details on these positions, detailed topics, and other information, contact Daniel Schaerer

(schaerer@ast.obs-mip.fr).
Related information also available from the URL http://webast.ast.obs-mip.fr/people/schaerer/